How do I know which one to subtract by?
EX. -7x-3y=42
Would I subtract by -7x or -3y?

Answer by tutorcecilia(2152)

-7x-3y=42 [You are trying to find the value of y, so isolate the y term. Because you are trying to re-write the equation in the slope-intercept form of (y=mx+b), so in this case, isolate the y]
.
-7x+7x-3y=+7x+42 [Add (7x) to both sides]
-3y=7x+42 [Divide both sides by (-3)]
-3y/-3=7x/-3+42/-3
y=(-7/3)x+(-42/3) [Simplify]
y=-7/3x-14